Solve the system.
x – 4y – z = 21
6x – 3y – z = –4
–x +2y –5z = 19

Answer by Alan3354(69443)

Solve the system.
x – 4y – z = 21
6x – 3y – z = –4
–x +2y –5z = 19
--------------------
Step 1, Subtract the 2nd eqn from the 1st.
Step 2, multiply the 1st eqn by 5, then subtract from the 3rd.
----
z is gone.